Circular No. 6296

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1996 B1 Edward W. Szczepanski, Houston, TX, reports his discovery of a

comet, observations of which are available as follows:

1996 UT  α_{2000}  δ_{2000}  m_{1} Observer
Jan. 27.455 14^{h}08.^{m}0 +53^{o}54′ 10.5 Szczepanski
28.33 14 05.2 +53 44 9.3 Drake
28.375 14 04.9 +53 43  Mitchell
28.415 14 04.8 +53 45  Hale
28.449 14 04.6 +53 45 8.2 _{”}

E. W. Szczepanski (Houston Astronomical Society). 0.10-m refractor and 300-mm camera lens. 50-min exposures show diffuse, trailed images similar in appearance to NGC 5474, 30′ to southwest.

K. Drake (Houston Astronomical Society). 0.44-m reflector. Coma 5′.5 × 4′

in p.a. 80^{o}, condensation shifted slightly west of center. L. Mitchell (Houston Astronomical Society). 0.3-m reflector. Communi-

cated by B. Wilson. A. Hale (Cloudcroft, NM). 0.41-m reflector (position) and 10×50 binoculars

(magnitude). Coma diameter 11′, prominent central brightening slightly offset toward leading side of coma.

CATALOGUE OF COMETARY ORBITS 1996 A new edition of the Catalogue of Cometary Orbits, complete through

the end of 1995, is now available. It utilizes the new system of cometary designations but includes detailed cross-references with old systems. Separate entries are given for individual components of several split comets, yielding a total of 1498 entries for 1470 cometary apparitions referring to 883 different comets, 185 of them of short period. There is a tabulation giving osculating elements for the 124 numbered periodic comets (except for five that are lost) for the standard epochs 1996 Apr. 27 and Nov. 13. ‘Original’ and ‘future’ reciprocal semimajor axes are given for the 291 highquality long-period orbits. The 110-page Catalogue sells for $20.00 ($30.00 for airmail delivery outside North America). There are also diskette and e-mail editions with the basic orbital information (the former also with a program for computing ephemerides), and these cost $110.00 ($120.00 by airmail) and $50.00, respectively.

1996 January 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden